Suppose that many stocks are traded in the market and that it is possible to borrow at the risk-free rate, rƒ. The characteristics of two of the stocks are as follows:

Stock Expected Return Standard Deviation
A 6 % 20 %
B 10 % 80 %
Correlation = –1


a. Calculate the expected rate of return on this risk-free portfolio? (Hint: Can a particular stock portfolio be substituted for the risk-free asset?) (Round your answer to 2 decimal places.)
Answer #1

Since the stocks are completely negatively correlated, the portfolio variance is zero.

wA = weighting of asset A
wB = weighting of asset B
sdA = standard deviation of asset A
sdB = standard deviation of asset B
p = correlation of A and B

portfolio variance = [wA2 * sdA2] + [wB2 * sdB2] + [2 * wA * wB * p * sdA * sdB]

Here p=-1 and implies wAsdA = wBsdB

Let wB be (1-wA)

0.20* wA = 0.80 * (1 - wA)

0.20* wA = 0.80 - (0.80 * wA)

[0.20*wA] + [0.80 * wA] = 0.80

wA = 0.80, or 80%

wB = 1 - wA = 1 - 0.80 = 0.20, or 20%

E(R) = [wA * rA] + [wB * rB]

= (0.80 * 6%) + (0.20 * 10%)

= 4.80% + 2% = 6.80%